OpenStudy (anonymous):

square root of 72 divided by the square root of 2

OpenStudy (anonymous):



OpenStudy (saifoo.khan):

\[\frac{\sqrt{72}}{\sqrt2} \times \frac{\sqrt2}{\sqrt2}\]

OpenStudy (saifoo.khan):

Can u solve now?

OpenStudy (anonymous):

keep going please!

OpenStudy (paxpolaris):

\[\large {\sqrt {72} \over \sqrt2}= \sqrt {72 \over 2}\]

OpenStudy (saifoo.khan):

That's a nice way too^

OpenStudy (anonymous):

wait shouldnt it be 72 x 2 ______ 2

OpenStudy (anonymous):

since u multiply both by \[\sqrt{2}\]

OpenStudy (saifoo.khan):

My method = \[\frac{\sqrt{72} \times \sqrt2 }{2}\]

OpenStudy (anonymous):

okay, whats the next step...

OpenStudy (saifoo.khan):

\[\frac{\sqrt{72 \times 2}}{2}\]

OpenStudy (saifoo.khan):

\[\frac{\sqrt{144}}{2} \to \frac{12}{2} \to 6\]

OpenStudy (paxpolaris):

or for this question you could do \[\large {\sqrt {72} \over \sqrt2}= \sqrt {72 \over 2}= \sqrt{36}=\Large 6\]

OpenStudy (anonymous):